Grade 10Physical Chemistry

Find the equivalent weight of H3PO4 in the reaction:
Ca(OH)2 + H3PO4 = CaHPO4 + 2H2O

Hrishant Goswami
12 Years agoGrade 10

2 Answers

Jitender Pal
12 Years ago
∵ In the given reaction two hydrogen atoms of H3PO4 are replaced so the basicity of H2PO4 is 2.
Using the relation, eq. wt. of acid = Mol. wt. of acid/Basicity
We get, eq. wt. of H3PO4 = Molwt/2 = (1 *3) +31 +(4 * 16)/2
= 98/2 = g
Deepak Patra
12 Years ago
λmin = hc / eV = 6.63 x 10-34 x 3 x 108 / 1 . 6 x 10-19 x 20 x 103 = 0.62 Å